if x, y, and z are the first three terms of a geometric sequence, show that x^2, y^2 and z^2 form another geometric sequence
Answer by mananth(16946)

x y z are in geometric
x/y = y/z
OR y^2=xz

x^2, y^2, z^2
x^2, xz, z^2
If they are in geometric sequence then
xz/x^2 = z^2/xz
z/x = z/x
The ratios are common
so x^2,y^2 & z^2 is a GP